You are given a 2.9265 g sample of a substance with the generic formula MCl2(H2O)6. After completely drying the sample (which means removing the 6 moles of H2O per mole of MCl2), the sample weighs 1.5969 g.
Answer #1

mass of H₂O = 9.9265-1.5969 = 1-32969 mass of Mug = 1.59699 mass of smoles of H₂O = 1089. be the half of Melo mass in 1089 of